# Context-Guided Spatial Feature Reconstruction for Efficient Semantic Segmentation

### v1

@@ 3.3 Dynamic Prototype Guided Head
To improve the classification of foreground objects, the Dynamic Prototype Guided head is proposed. The DPG head generates the dynamic prototype to embed class information. The class embedding feature is projected into the pixel feature space and weight pixel features to enhance the distinction between different classes. The overall architecture of DPG Head is shown in Fig. 4.
@@
Then, the DPG head further embedded class information to enhance the distinction between different classes. A fully connected layer is applied to compress the prototype to C×$\times$1 dimension to embed the class information and generate the class embedding vector, and Softmax is utilized to constrain the value between 0 and 1. The class embedding vector 1$\times$×C represents the global information in each class. To project the class embedding vector into the pixel feature space, we multiply the class embedding vector by the transposed prototype. With the class embedding enhancement, the new attention vector has a stronger ability to distinguish the classes. Then, two fully connected layers are used to capture the context in the attention vector. Layer Normalization is utilized to normalize features between two layers. The calculation process of the class embedding vector is summarized by:
$\rm\mathbf{F}_{gp}=Softmax\left(\delta_{D\rightarrow 1}\left(\mathbf{F}_{p}\right)\right)$ (6)
@@ 4.1 Datasets
ADE20K: The ADE20K [38] is a challenging segmentation dataset. The dataset contains more than 20,000 images from diverse environments and situations, covering 150 categories. All images are exhaustively annotated with pixel-level objects and object parts labels. It consists of 20,210/2,000/3,352 images in the training, validation, and test sets.
COCO-Stuff: COCO-Stuff is built from COCO dataset by augmenting all 164K images with pixel-level stuff annotation. It contains $172$ categories with $80$ things, $91$ stuff and $1$ unlabeled class. There are 118K/5K/20K/20K images for training, validation, test-dev and test-challenging.
Pascal Context: The PASCAL Context dataset is an extension of the PASCAL VOC 2010 detection challenge, which contains 59 classes. The dataset contains 10100 images, including 4996 images in the training set and 5104 images in the test set.
Table 1: Results on ADE20K. We report the performance in terms of the mIoU, FLOPs, Param, and Throughputs. The top rows, middle rows and bottom rows are tiny models, base models and large models, respectively.
@@
Mask2Former(CVPR’22) [8] 47.7 74.0 63.0 -
@@ 4.2 Implementation details
Our implementation is based on the public codebase mmsegmentation [10]. We use EfficientFormerV2 [20] as the backbone network. For the ADE20K dataset, we follow SegNeXt [14] to use 160K scheduler and the batch size is 16. The initial learning rate is 0.00012 and the weight decay is 0.01. A “poly” learning rate scheduled with factor 1.0 is adopted. The training images are resized to 1024$\times$512 for ADE20K. For the COCO-Stuff dataset, we use 80K scheduler and the batch size is 8. The initial learning rate is 0.00012 and the weight decay is 0.01. The training images are resized to 1024$\times$512. For all three datasets, we use the default data augmentation strategy in mmsegmentation. For the Pascal Context dataset, the training iteration is also 80K with batch size 16. The initial learning rate is 0.00006 and the weight decay is 0.01. The training images are resized to 480$\times$480.
@@ 4.3 Comparisons with the State-of-the-art Methods
For the tiny models, the proposed CGRSeg-T achieves $43.6\%$ mIoU with only $4.0$ GFLOPs. It performs better than all other CNN-based models and transformer-based models in terms of the mIoU and FLOPs. For example, it exceeds the previous best model SeaFormer-L by $+0.9\%$ mIoU with $2.5$ GFLOPs lower computation cost and $4.6$M lower parameters. As for the base models, the proposed CGRSeg-B yields $45.5\%$ mIoU with $7.6$ GFLOPs. Note the computation cost of our model is only half of the best compared SegNeXt-S ($15.9$ GFLOPs) model, but the performance is $+1.2\%$ mIoU better. Among the large models, our CGRSeg-L model ($15.1$ GFLOPs) outperforms MaskFormer (Swin-T) and Mask2Former (Swin-T) over $+1.6\%$/$+0.6\%$ mIoU, but only with around $1/4$ computational cost, which demonstrates the high efficiency of our model.
Based on the above experimental results, we can conclude that the proposed CGRSeg can achieve higher performance with less computation cost under the same experimental setup. This is because the proposed RCM can make the model focus more on the key areas and improve the feature representation. Therefore, our model can achieve high segmentation accuracy at a small computation cost.
@@
Results on Pascal Context. We compare the proposed CGRSeg with state-of-the-art models on the Pascal Context dataset. As shown in Table 3, our approach achieves better performance than all other compared methods. For example, our CGRSeg-T model obtains $54.1\%$ mIOU with only $4.0$ GFLOPs, which is $+0.8\%$ better than SegNeXt-T [14] with $6.6\%$ FLOPs. When compared with the larger model SegNeXt-S, our model still achieves $+0.4\%$ better mIoU with about only $48\%$ FLOPs. Besides, CGRSeg-B also exceeds Senformer [1] $2.2\%$ mIoU.
@@ 4.4 Ablation Study
The ablation analysis of the proposed modules. The two core components of the proposed CGRSeg are the RCM and the final DPG head. To validate the effectiveness of the two components, in Table 4, we make an ablation of the RCM and DPG head. The base model is the first row of the table without any elaborate designs. The proposed RCM is used for pyramid context extraction (PCE) and spatial feature reconstruction (SFR). In base model, we replace the RCM for pyramid context extraction by a $1\times 1$ convolution, and directly drop RCMs for feature reconstruction. DPG Head is directly dropped. As shown in the first row of the table, when not using the two modules, the performance is only $40.86\%$ mIoU. In third row, adding the RCM for pyramid context extraction can increase the performance by $1.23\%$ mIoU with only an extra $0.19$ GFLOPs. Adding the RCM for feature reconstruction can increase the performance by $1.13\%$ mIoU with only an extra $0.17$ GFLOPs. Adding the DPG Head, i.e., the second row in the table, the performance will reach $41.34\%$ from $40.86\%$. When combined with the two modules, the final performance can achieve $43.60$, which is $2.74\%$ mIoU than the base model, i.e., the first row of the table.
@@
None-local 40.9 4.05 9.4
DAHead 42.6 4.05 9.4
HamHead 42.1 4.05 9.4
GC block 42.7 3.92 9.3
SE block 42.6 3.92 9.4
DPG Head 43.6 4.00 9.4
Table 6: The performance comparison of RCA and other advanced attention block. RCA significantly outperforms other attention modules without increasing Flops.
@@
Comparisons of alternative blocks of the RCA. As shown in Table 6, we compare the performance of other alternative sub-blocks of the RCA, e.g., self-attention [35], InceptionNext [30], ConvNext [22],oordAtt Block [15] and GatherExcite [16] Block. To compare them, we directly replace the RCM with other blocks. It can be found that the proposed CGRSeg performs better than the CoordAtt Block and GatherExcite Block by increasing $2.1\%$ and $1.9\%$ mIoU.
@@
The effectiveness of the kernel size in shape self-calibration function. To investigate the effectiveness of the kernel size in the shape self-calibration function, we conduct a series of experiments in the Table 10. As the size of the convolution kernel increases, the segmentation precision also increases. When the kernel sizes are $1\times 1$ 11, and $11\times 1$, the performance is best. Because increasing the kernel size further leads to the increase of FLOPs, we do not increase the convolution kernel. Thus, we finally choose the groups of $1\times 1$ 11, and $11\times 1$ kernel sizes in our model.
Convolution kernel selection for fusion functions in RCA. To investigate the effectiveness of the kernel size in the fusion functions, we conduct a series of experiments in the Table 10. As shown in the table, when the kernel sizes are $3\times 3$, the performance is best. Thus, we finally choose $3\times 3$ kernel sizes in fusion function.
@@
Visualization. The prediction results of SegNeXt and CGRSeg-T are visualized in Figure 5. The mask of CGRSeg is closer to the ground truth than that of other methods. It can be found that our CGRSeg-T achieves more complete segmentation masks with fewer errors. The mask is more complete because RCM can make the model pay more attention to the foreground. Besides, the DPG head can embed with category information can reduce classification errors.
@@ 5 Conclusion
In this paper, we present CGRSeg, an efficient yet competitive segmentation framework which consists of pyramid context extraction and spatial feature reconstruction. To implement these two parts, a RCM is carefully designed to make model focus on the target area and capture axial global context. Also, the DPG Head is proposed to significantly improve the feature representation with only a small amount of computation. Our CGRSeg is extensively evaluated on ADE20K, Pascal Context, and COCO-Stuff benchmarks, and achieves the state-of-the-art segmentation performance on all three datasets.

### v2

@@ 3.3 Dynamic Prototype Guided Head
To improve the classification of foreground objects, the Dynamic Prototype Guided head is proposed, which generates the dynamic prototype to embed class information. The class embedding feature is projected into the pixel feature space, weighting the pixel features to enhance the distinction between different classes. The overall architecture of DPG Head is shown in Fig. 4.
@@
Then, the DPG head further embeds class information to enhance the distinction between different classes. A fully connected layer is applied to compress the prototype to C×$\times$1 dimension, effectively embedding the class information and yielding a class embedding vector. Softmax is applied to constrain the value between 0 and 1. The class embedding vector C×$\times$1 represents the global information in each class. To project the class embedding vector into the pixel feature space, we multiply the class embedding vector by the transposed prototype. With the class embedding enhancement, the new attention vector has a stronger ability to distinguish the classes. Then, two fully connected layers are used to capture the context in the attention vector. Layer Normalization is utilized to normalize features between two layers. The calculation process of the class embedding vector is summarized by:
$\rm\mathbf{F}_{gp}=Softmax\left(\delta_{D\rightarrow 1}\left(\mathbf{F}_{p}\right)\right),$ (6)
@@ 4.1 Datasets
We conduct comprehensive experiments on the ADE20K [45], COCO-Stuff [2], and Pascal Context [28] datasets. The ADE20K [45] is a challenging segmentation dataset. It contains more than 20,000 images from diverse environments and situations, covering 150 categories. COCO-Stuff [2] is built from the COCO dataset by augmenting 164K images with pixel-level stuff annotation. It contains $172$ categories with $80$ things, $91$ stuff, and $1$ unlabeled class. The PASCAL Context [28] dataset is an extension of the PASCAL VOC 2010 detection challenge, which contains 59 classes. The dataset contains 4996 images in the training set and 5104 images in the test set.
Table 1: Results on ADE20K. We report the performance in terms of the mIoU, FLOPs, Parameter, and Throughput.
@@
Mask2Former(CVPR’22) [10] 47.7 74.0 47.0 -
@@ 4.2 Implementation details
Our implementation is based on the public codebase mmsegmentation [12]. We use EfficientFormerV2 [23] as the backbone network. For the ADE20K dataset, we follow SegNeXt [16] to use 160K scheduler and the batch size is 16. The initial learning rate is 0.00012 and the weight decay is 0.01. A “poly” learning rate scheduled with factor 1.0 is adopted. The training images are resized to 1024$\times$512 for ADE20K. For the COCO-Stuff dataset, we use 80K scheduler and the batch size is 8. The initial learning rate is 0.00012 and the weight decay is 0.01. The training images are resized to 1024$\times$512. For the Pascal Context dataset, the training iteration is also 80K with batch size 16. The initial learning rate is 0.00006 and the weight decay is 0.01. The training images are resized to 480$\times$480. For all three datasets, we use the default data augmentation strategy in mmsegmentation. The throughput is evaluated on an NVIDIA Tesla V100 GPU with the maximum batch size.
@@ 4.3 Comparisons with the State-of-the-art Methods
For the tiny models, the proposed CGRSeg-T achieves $43.6\%$ mIoU with only $4.0$ GFLOPs. It performs better than all other CNN-based models and transformer-based models in terms of the mIoU and FLOPs. For example, it exceeds the previous best model SeaFormer-L by $+0.9\%$ mIoU with $2.5$ GFLOPs lower computation cost and $4.6$M lower parameters. As for the base models, the proposed CGRSeg-B yields $45.5\%$ mIoU with $7.6$ GFLOPs. Note the computation cost of our model is only 39% of the best compared PEM-STDC2 ($19.3$ GFLOPs) model, but the performance is $+0.5\%$ mIoU better. Among the large models, our CGRSeg-L model ($15.1$ GFLOPs) outperforms MaskFormer (Swin-T) and Mask2Former (Swin-T) over $+1.6\%$/$+0.6\%$ mIoU, but only with around $1/4$ computational cost, which demonstrates the high efficiency of our model.
Based on the above experimental results, we can conclude that the proposed CGRSeg can achieve higher performance with less computation cost. This is because the proposed lightweight RCM can make the model focus on the foreground objects, improving the feature representation. Additionally, the DPG Head also improves the classification of the foreground objects.
@@
Results on Pascal Context. We compare the proposed CGRSeg with state-of-the-art models on the Pascal Context dataset. As shown in Table 3, our CGRSeg achieves better performance than all other compared methods. For example, our CGRSeg-T model obtains $54.1\%$ mIOU with only $4.0$ GFLOPs, which is $+0.8\%$ better than SegNeXt-T [16] with $6.6$ FLOPs. When compared with the larger model SegNeXt-S, our model still achieves $+0.4\%$ better mIoU with about only $48\%$ FLOPs. Besides, CGRSeg-B also exceeds Senformer [1] $2.2\%$ mIoU.
@@ 4.4 Ablation Study
The ablation analysis of the proposed modules. The two core components of the proposed CGRSeg are the RCM and the final DPG head. To validate the effectiveness of the two components, in Table 4, we make an ablation study of the RCM and DPG head. The base model is the first row of the table without any elaborate designs. The proposed RCM is used for pyramid context extraction (PCE) and spatial feature reconstruction (SFR). In the base model, we replace the RCM for pyramid context extraction by $1\times 1$ convolutions and directly drop RCMs for feature reconstruction. DPG Head is directly dropped. As shown in the first row of the table, when not using the two modules, the performance is only $40.86\%$ mIoU. In the third row, adding the RCM for pyramid context extraction can increase the performance by $1.23\%$ mIoU with only an extra $0.19$ GFLOPs. Adding the RCM for feature reconstruction can increase the performance by $1.13\%$ mIoU with only an extra $0.17$ GFLOPs. Adding the DPG Head, i.e., the second row in the table, the performance will reach $41.34\%$ from $40.86\%$. When combined with the two modules, the final performance can achieve $43.60\%$ mIoU, achieving a $2.74\%$ improvement over the base model.
@@
None-local 40.9 4.05
DAHead 42.6 4.05
HamHead 42.1 4.05
GC block 42.7 3.92
SE block 42.6 3.92
DPG Head 43.6 4.00
Table 6: The performance comparison of RCA and other advanced attention block. RCA outperforms other attention modules without increasing Flops.
@@
Comparisons of alternative blocks of the RCA. As shown in Table 6, we compare the performance of other alternative sub-blocks of the RCA, e.g., self-attention [42], InceptionNext [37], ConvNext [26], CoordAtt Block [17] and GatherExcite Block [18]. To compare them, we directly replace the RCM with other blocks. It can be found that the proposed CGRSeg performs better than the CoordAtt Block and GatherExcite Block by increasing $2.1\%$ and $1.9\%$ mIoU.
@@
The effectiveness of the kernel size in shape self-calibration function. To investigate the effectiveness of the kernel size in the shape self-calibration function, we conduct a series of experiments in the Table 10. As the size of the convolution kernel increases, the segmentation precision also increases. When the kernel sizes are $1\times 1$ 11, and $11\times 1$, the performance is best. Because increasing the kernel size further leads to the increase of FLOPs, we do not increase the convolution kernel. Finally, we choose the groups of $1\times 1$ 11, and $11\times 1$ kernel sizes in our model.
Convolution kernel selection for fusion functions in RCA. To investigate the effectiveness of the kernel size in the fusion functions, we conduct a series of experiments in the Table 10. As shown in the table, when the kernel size is $3\times 3$, the performance is best. Thus, we finally choose $3\times 3$ kernel size in the fusion function.
@@
Visualization. The prediction results of SegNeXt and CGRSeg-T are visualized in Fig. 5. The mask of CGRSeg is closer to the ground truth than that of other methods. It can be found that our CGRSeg-T achieves more complete segmentation masks with fewer errors.
@@ 5 Conclusion
In this paper, we present CGRSeg, an efficient yet competitive segmentation framework consisting of pyramid context extraction and spatial feature reconstruction. To implement these two parts, an RCM is carefully designed to make the model focus on the foreground and capture the axial global context. Furthermore, the DPG Head is proposed to significantly improve the feature representation with only a small amount of computation. Our CGRSeg is extensively evaluated on ADE20K, COCO-Stuff, and Pascal Context benchmarks, achieving state-of-the-art segmentation performance on all three datasets.